Subject: Minutes of April 8, 1996 User's Committee Meeting

From: Michael Strauss

Submitted: Mon, 15 Apr 1996 08:52:01 -0400 (EDT)

Message number: 58 (previous: 57, next: 59 up: Index)

  Minutes of APO 3.5m User's Committee Phone Conference
		April 8, 1996

  Attending: Alan Uomoto, Bruce Gillespie, Ed Turner (Chair), Lew
Hobbs, Roger Hildebrand, Michael Strauss (taking minutes), Rene
Walterbos, Alan Uomoto, Julie Lutz, Chris Stubbs

  Current status of Echelle
  Plans for realuminizing the primary
  Optical throughput of telescope/instruments
  Recent engineering projects done by UWashington and APO people
  Preliminary thoughts for a JHU/UC/ARC instrument

 What is the current status and installation plan for the Echelle?

Hildebrand: Last summer ran into a serious vacuum problem, leading to
major redesign. New camera tank being fabricated, ready early May; 6
weeks of testing of this tank, then mount CCD in camera (will relevant
manpower be busy with Sloan camera?). Six weeks of further testing.
Following this schedule, the Echelle could be delivered in October.
Now at high confidence that this *will* be completed successfully.
Just about everything with the instrument is being rebuilt from
  One small issue: The Chaos laser is currently sitting at the port at
which the echelle will go; the Echelle and Chaos people need to
coordinate this. 
  It is not yet clear how much time commissioning will take. One
potential problem is that Bob Lowenstein, whose expertise will be
critical during commissioning, is planning a several-month trip to
the South Pole around October, which could delay things several

  With this in mind, the UC suggests that proposals not be accepted
for survey instruments until the quarter *following* the declaration
that they are successfully commissioned. 

  We also recommended that the commissioning time for this instrument
would come out of engineering time (i.e., it would be shared between
institutions, and not just come out of U Chicago time). 

  The HRI (High-Resolution Imager) is not currently being worked on;
there are no plans at the moment to reserrect it. 

  Plans for Coating the Mirrors
  Given the fear (see below) that the mirror coatings are causing
significant loss of reflectivity, especially in the UV, we are looking
to do recoating of the primary in late June at KPNO (our next
opportunity would be October). There may be a conflict with the need
to get the 2.5m mirror coated; if so, the 2.5m mirror could perhaps be
done at McDonald.  Decision will be made early May, depending on the
details of McDonald's setup.  This will of course have an effect on
the telescope schedule.  Secondary and tertiary coatings are bare
aluminum.  More sophisticated coatings have better reflectivity over
large wavelength range (down to 3500 A) and are easy to clean.  We
might look into this.

 Optical throughput of the telescope/instruments

Tonight (4/8), there will be a run measuring throughput of DIS in some
detail (see telescope logs for the results of this run), to be carried
out by NMSU personell. This is engineering time.  Will follow Jim
Gunn's suggestions for tests of throughput.  Tests of slit masks as

  Jim Gunn suspects that the throughput problems are related to mirror
coatings, especially in blue.  Observatory reflectometer only works in
red.  Multiwavelength devices don't exist.  Jim will design a device
to do this; APO is offering to build it.  
  Spectrophotometric calibration device now is at APO for measuring
throughput of MT; use it for 3.5m as well! 

  We did not get a chance to discuss options/plans for new slits for

Engineer projects--recent work done by the "Improvements" group at
          UWash, APO.

  Bruce Gillespie: Last week of engineering time: new, stronger
rotator motor has been installed (this motor has overcurrented;
problems are not over yet).  Tightening of secondary vanes has changed
the focus as function of altitude has changed; data were taken to
recalibrate this, and the run of focus as a function of
temperature. This should allow a better automatic focus adjustment.
Ideally, we could have a semi-automatic mechanism by using the guider
to continually measure focus (and seeing, of course), and update
pointing model as well.
  Chris Stubbs: Telescope vibrations: instability in telescope servo
systems and wind-driven oscillations are important contributors to
seeing. Secondary vane tightening seems to have improved the
oscillation problems. Fixes to servo problems are in progress.  There
is some evidence that the 20 Hz vibrational problem is much reduced.
  Progress on reducing the effects of scattered light is going slow,
due to lack of manpower. 
  The worst of the enclosure wheels was successfully fixed. It would
take a 7-10 day shutdown to fix the other three wheels (perhaps this
could be done during realuminization in late June?).
  How do we know that secondary is really at fault in optics?  The
best way to do this is to do an optical test of secondary; take it off
the telescope and ask someone to map it. The test of rotating the
secondary 120 degrees would be inconclusive, given the inherent noise
in the Hartman tests. We could do this while primary is recoated in
late June.

  Preliminary thoughts for a JHU/UC/ARC instrument

  Alan Uomoto summarizes discussions at JHU about this instrument:
instrument should be capable of unique science, must have JHU
interest.  Should it be a key project kind of instrument?  Current
favorite idea is a wide-field (12-15') low dispersion (R = 1000) high
throughput multiobject (40) spectrograph. Using aperture masks, not
fibers. Larger field of view than instruments on Gemini, Keck.
Wavelength coverage 4000-7500 A in one shot. Think of this as a Sloan
followup instrument. Can use Advanced Camera CCD's (2048x4096). This
is one of a few options still being discussed. Chicago is interested
in add-on instruments for Chaos. Time to get input from whole APO
community; Alan will put together a several-page description of their
idea for comments from everybody.

  Finally, it was decided that the APO community should have the
opportunity to comment on the minutes after they are posted, to
correct any technical mis-statements. People should send corrections
and/or comments to Michael Strauss ( or 
to apo35-general, and the appropriately corrected minutes will be
approved at the start of each User's Committee meeting.

  Next meeting, May 13, at usual time. 

Two items for the agenda next time:
        o data archiving at APO--is a minimal semi-automated archive of image
          data at the site worth the cost and added work?  Something on the
          scale of the "Save the Bits" program at KPNO.
	 Recently, people have lost data by improper ftp'ing. 

        o Network and remote interface problems and enhancement plans.
	The NSF has recently put out a call for proposals for
	netwrk/throughput-related issues. Don York may already be on

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 58 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to